Aqueous solution of HCL is diluted by fuve times. What will be thr change in pH? 

Grade:12

1 Answers

Naveen Kumar
9 years ago
As we know, pH=-log[H+]
let initailly, [H+]=x, So pH=-log[H+]=-logx
after diluting the acid 5 times, the concentration of [H+]=x/5, Hence pH=-log[H+]
.............................................................................=-log(x/5)=-logx+log5=-logx+0.7
So change in pH=0.7